Please help me solve these 2 problems:
8 times the difference of a number and 3 is 16.
Solve 3xyz-2xy+3xz-4x+5y for x=-3, y=-3, and z=-4.

Answer by xcentaur(357)

Let the number be x
8(x-3)=16
8x-24=16
8x=40
x=5


3xyz-2xy+3xz-4x+5y for x=-3, y=-3, and z=-4


3(-3)yz-2(-3)y+3(-3)z-4(-3)+5y [Substituting x]
-9yz+6y-9z+12+5y
-9(-3)z+6(-3)-9z+12+5(-3) [Substituting y]
27z-18-9z+12-15
27z-9z-6+15
18z+9
18(-4)+9 [substituting z]
-72+9
-63


Hope this helps,
Prabhat